OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. (more details to come)

OpenStudy (anonymous):

\[f(x)=(x-7)/(x+3) and g(x)=(-3x-7)/(x-1)\]

OpenStudy (anonymous):

@hero Any ideas? I'm not really sure how to go about this... unless you're supposed to turn it into a complex fraction?

hero (hero):

\[f(x) = \frac{x - 7}{x + 3}\] \[g(x) = -\frac{3x+7}{x - 1}\]

hero (hero):

Notice that each fractional expression is improper. What you do is re-write each fraction as a mixed fraction.

hero (hero):

When you do this, the resulting expression will have only one x variable. This is will make it easier for you to do f(g(x)) and g(f(x)).

hero (hero):

To convert f(x) to a mixed fraction: \(f(x)=\dfrac{x - 7}{x + 3} \) \(= \dfrac{x + 3 - 10}{x + 3}\) \(=\dfrac{x + 3}{x + 3} - \dfrac{10}{x + 3}\) \(=1 - \dfrac{10}{x + 3}\)

hero (hero):

Notice that f(x) is now expressed as a mixed fraction with only one \(x\) variable as input

hero (hero):

Now to convert g(x) to a mixed fraction, we will do a similar process:

hero (hero):

\[g(x) = -\frac{3x+7}{x - 1}\] \(= -\dfrac{3x - 3 + 10}{x - 1}\) \(=-\dfrac{3x - 3}{x - 1} - \dfrac{10}{x - 1}\) \(=-3 - \dfrac{10}{x - 1}\)

hero (hero):

So now, \(f(g(x)) = 1 - \dfrac{10}{-3-\frac{10}{x-1} + 3}\) \(= 1 - \dfrac{10}{-3 + 3-\dfrac{10}{x-1}}\) \(=1 - \dfrac{10}{-\dfrac{10}{x-1}}\) \(=1 + \dfrac{10}{\dfrac{10}{x-1}}\) \(=1 + 10 \times \dfrac{x - 1}{10}\) \(=1 + x - 1\) \(=1 - 1 + x\) \(=x\)

hero (hero):

I hope you are able to understand how I got that.

hero (hero):

Do a similar process for g(f(x)) to arrive at x

OpenStudy (anonymous):

thank you!

hero (hero):

Did you understand any of that?

hero (hero):

If you have any questions, now is the time to ask. I won't be around much longer.

OpenStudy (anonymous):

Yeah, I ended up with the correct answer.

hero (hero):

Did you calculate g(f(x)) yet?

hero (hero):

@Precalcstuent2013, it is possible that you may be having trouble with this. I skipped some steps just for the sake of posting a solution in a timely manner. It is easy to confuse some of the steps. So if you are having trouble with it, please let me know so I can clarify things. There is no need to go on being frustrated about any of the steps if you are having trouble. The method I have shown you is not a popular one.